### v1

# Planning from Imagination: Episodic Simulation and Episodic Memory for Vision-and-Language Navigation

Abstract
Humans navigate unfamiliar environments using the capabilities of episodic simulation and episodic memory. Developing imagination-based memory, analogous to episodic simulation and episodic memory, can enhance embodied agents’ comprehension of the complex relationship between environments and objects. However, existing Vision-and-Language Navigation (VLN) agents fail to perform the aforementioned mechanism. We propose a novel architecture to help agents build a recurrent imaginative memory system. Specifically, the agent can maintain a reality-imagination hybrid global memory during navigation and expand the memory map through imaginative mechanisms and navigation actions. Correspondingly, we design a series of pre-training tasks to help the agent acquire fine-grained imaginative abilities. Our agents improve the state-of-the-art (SoTA) success rate (SR) by 7% while simultaneously imagining high-fidelity RGB representations for future scenes.
1 Introduction
Autonomous embodied agent navigation is advancing through Vision-and-Language Navigation (VLN) research (Anderson et al. 2018; Qi et al. 2020). In a VLN task, the agent needs to reach the target location given navigation instructions (e.g. “Exit the room then go straight and turn left. Go straight until you pass a TV.”). This task, however, is more challenging when agents navigate in unseen environments. The navigation performance is significantly degraded compared to that in the seen environment.
Thus, it is natural to focus on human navigation behaviors in unseen environments, and two important neuroscience concepts were identified: the episodic simulation and the episodic memory (Kühn and Gallinat 2014; Gomez, Rousset, and Baciu 2009), which means using memories to imagine uncertain information or scenes, aiding future decision-making mentally). According to their navigation memory, humans can imagine both fine-grained knowledge (e.g., RGB images) and higher-level spatial information (e.g., location distribution) for the future in unseen environments (Tang, Yan, and Tan 2017). For example, if the navigation instruction is “go to the laundry room” (Fig 1.), humans would imagine a room with washing machines, likely to be adjacent to a toilet they passed earlier. Such a simulation would provide clues for navigating directions.
Existing approaches equip agents with mechanisms to infer or imagine unseen environments. However, these agents lack the ability to integrate their imaginative outputs into long-term memory, preventing the formation of episodic simulations and persistent memories (Wang et al. 2023a; An et al. 2023). While current methods enable agents to predict potential object locations (Shah et al. 2023; An et al. 2024) and RGB features (Moudgil et al. 2021; Qi et al. 2021), such predictions are often ephemeral and transient, overwritten at each navigation step. This transient nature prevents agents from constructing durable memory representations that fuse visual and contextual information to support future reasoning or decision-making. For instance, in Fig 1, the inability to infer the presence of spice bottles from contextual cues such as a knife and fork highlights the limitations of existing imagination mechanisms. Consequently, these methods produce isolated and short-lived imaginative outputs, falling short of the continuity and compositional depth characteristic of human episodic memory and simulation.
Figure 1: Humans utilize episodic memory and episodic simulation to recall past experiences and predict future outcomes in unfamiliar environments. However, navigation agents often struggle in unseen environments due to their inability to construct and leverage such cognitive frameworks.
In this paper, we introduce the Space-Aware Long-term Imaginer (SALI) agent, the first agent designed to emulate human-like episodic simulation and memory for navigation. The SALI agent leverages an imagination-based memory system to capture both high-level spatial information about the navigation environment and fine-grained local features. By incorporating commonsense knowledge (e.g., associating specific objects with corresponding rooms, such as a refrigerator with a kitchen), the SALI agent builds and maintains a long-term imagination-based memory, as illustrated in Fig 2. This memory structure enables the agent to integrate imaginative outputs with stored information to support informed action decisions. To implement this capability, we model the agent’s memory using a topological map structure (Chen et al. 2022), where nodes represent spatial locations and are enriched with local RGBD data and semantic information generated by the imagination component. Multimodal transformers are employed to encode these nodes alongside natural language instructions, enabling the agent to make effective navigation decisions at each step. The combination of imagination and memory ensures that the SALI agent can reason holistically about its environment while adapting dynamically to new scenarios. As a result, the SALI agent achieves state-of-the-art performance on two widely recognized VLN benchmarks, demonstrating the effectiveness of our approach.
To summarize, our contributions are as follows.
  • Our proposed SALI agent has human-like episodic memory and episodic simulation abilities, enhancing navigation performance in unseen environments. Viewing from the perspective of human brains, we endow the agent with the ability to learn commonsense knowledge through imagination and memory.
  • We established a recurrent and end-to-end imagination mechanism that generates high-fidelity RGB scene representations. The SALI agent is capable of imagining realistic scene images and dynamically fusing the imagination into a real-imaginary combined memory. The agent can make navigation decisions based on this hybrid map, developing more effective navigation skills.
  • The SALI agent achieves state-of-the-art (SoTA) performance on VLN benchmarks. In the unseen environments, it improves 5% SR on the R2R dataset and 7% on the REVERIE dataset.
2 Related Works
Vision-and-language Navigation. Using visual inputs and natural language to guide robot navigation is a fundamental task in embodied AI (Anderson et al. 2018; Qi et al. 2020), where effectively aligning multimodal visual-linguistic inputs remains a core challenge (Lin et al. 2022). Existing VLN approaches (Lu et al. 2019; Qi et al. 2021; Hong et al. 2021) often struggle in unseen environments, where agents fail to correlate visual observations, particularly unfamiliar materials and textures, with navigation instructions, resulting in degraded performance. To address this limitation, we propose a novel VLN architecture inspired by human episodic simulation and memory, enabling agents to enhance navigation robustness and performance in unseen environments.
Memory Mechanism. Constructing memory to provide long-term decision support is important for embodied agents. SLAM has been widely used to construct memory maps, especially in navigation (Chaplot et al. 2020; Temeltas and Kayak 2008). To reduce memory overhead and enable more efficiency, the construction of memory mechanisms based on 2D bird-eye-view maps (An et al. 2023; Liu et al. 2023) or ego-centric grid maps (Georgakis et al. 2022; Wang et al. 2023c) can also help agents understand spatial information. Additionally, some navigation work uses topological maps to further simplify and extract high-level features to help the robot better learn spatial structure information (An et al. 2023). However, the correlation between historical memory and future imagination remains largely unexplored, leaving the potential of integrating the two underutilized. This oversight reduces the robustness of navigation agents and significantly impairs their performance in unseen environments. Meanwhile, the construction of abstract anticipatory and foresight thinking is necessary for the construction of human-like robots. We constructed a memory mechanism analogous to human episodic simulation memory to acquire knowledge about future scenarios.
Prediction during Navigation. Inspired by video prediction (Mallya et al. 2020), prior works have adopted imagery mechanisms to support navigation decisions, including predicting future spatial structures from depth camera data (Shah et al. 2023; Cui et al. 2019) and generating fine-grained visual information. They use RGB images to generate high-fidelity images (Moudgil et al. 2021; Wang et al. 2023a). Furthermore, there is also pioneering work that explores the generation of both spatial information and image information (Wang et al. 2018b; Li et al. 2023). Instead of simply applying imagination mechanisms in isolation, we leverage a memory-based framework to fully realize the potential of imaginative predictions. By integrating imagination into a memory mechanism, we enable the generation and maintenance of long-term, reusable navigation information. To achieve this, we propose an end-to-end adaptive image maintenance and generation mechanism that bridges the gap between transient predictions and persistent memory.
Figure 2: We propose building a hybrid imagination-reality memory for long-term navigation decisions. Based on the navigation observation and encoded features, the agent will generate a local imagination tree. This tree will then be fused into a hybrid memory, which the agent can utilize to make navigation decisions. The figure also illustrates a series of pre-training tasks that we propose for the imagery module.
3 Method
Problem Formulation. Our problem is set in a discrete VLN environment $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ (Anderson et al. 2018), where $\mathcal{V}$ represents navigation nodes and $\mathcal{E}$ represents the connectivity relations. The VLN dataset provides the instruction set $I$ and the corresponding ground-truth navigation paths $\hat{P}$. The VLN task requires guiding a robot to accomplish navigation tasks under $I$. At each navigation step $t$, the agent can acquire the RGB image $r_{t}$ and depth image $d_{t}$ via the camera and the current position information $p_{t}$ via the GPS sensor. Following (Irshad et al. 2022), agents are equipped with a visual classifier to acquire semantic images $s_{t}$. with observation $O_{t}=\{r_{t},d_{t},s_{t},p_{t}\}$, the agent needs to learn a policy $\pi(I,O_{t})$ to decide the current action $a_{t}$.
Method Overview. As shown in Fig 2, our SALI agent has a human-like episodic simulation and episodic memory mechanism. At each navigation step $t$, the agent will maintain a topological map as its memory to store both realistic and imagined information and make navigation decisions based solely on the map (Section 3.1). Then, the agent will use its memory to imagine future information of both high-level spatial knowledge and low-level image features and merge the imagery into the memory (Section 3.2). We conclude this part by presenting our approach to training and inference in Section 3.3.
Figure 3: The imagination model includes four pre-trained models (inpaint model, spade model, room-type model, and waypoint model). We propose an end-to-end architecture that allows local imagery trees to continuously update and maintain themselves, producing high-quality depth images, semantic images, and RGB images
3.1 Real-Imaginary Hybrid Memory
To provide the agent with the possibility of global action planning, we construct a mixed-granularity map memory structure. As shown in Fig 2, at navigation step $t$, the agent will update a topological map $G_{t}$ based on the current observation $O_{t}$ and historical information $H_{t}$.
3.1.1 Memory Map Representation
The global map can be represented as $G_{t}=\{N_{t},E_{t}\}$. Where the $E_{t}$ matrix records the Euclidean distance between neighboring nodes. $N_{t}$ represents the nodes in the map and records $V_{t}$, $p_{t}$ and $f_{t}$ information. $V_{t}=\left\{r_{t},d_{t},s_{t}\in\mathbb{R}^{H\times W\times C}\right\}$ in equirectangular panorama form will be used to generate the fine-grained images. Furthermore, we use the pre-trained Vision Transformer (ViT) (Lu et al. 2019) and ResNet encoder to get the image feature $f_{t}$, which is also stored in node $N_{t}$.
Stored nodes can be classified into four categories: visited nodes [Uncaptioned image], current nodes [Uncaptioned image], navigable nodes [Uncaptioned image], and imagination nodes [Uncaptioned image]. The information stored in these four types of nodes varies. For $V_{t}$, the complete and fixed image information has been stored in the visited node and current node. The navigable node’s image feature is represented by partial image features observed from neighboring real nodes and may contain information from imagination nodes. The $V_{t}$ of the imagination node is completely represented by the imagined image. For $p_{t}$, the agent cannot reach location $p_{t}$ of [Uncaptioned image], [Uncaptioned image], or [Uncaptioned image] via action $a_{t}$ (Anderson et al. 2018). Only the navigable node’s $p_{t}$ is the potential target location. In the map $G_{t}$, global action space is $A^{G_{t}}=\{a_{t}^{1},...a_{t}^{k}\}$, where $k$ is the number of navigable nodes in the map.
To reduce the memory load, we perform a pruning operation on the newly imagined nodes. The criteria of pruning are based on the sum of the feature cosine similarity and the negative position mean square error (MSE) between the imagination node $N_{i}$ and other navigable/imagination node $N_{j}$. If it is the same node, the node attributes are updated with an average pooling of the images and features from both nodes. The new node remains an imagination node when both are imagination nodes, otherwise is labeled as a navigable node. In addition, we set an upper bound of $\bar{N}$ on the number of imagination nodes to reduce memory load.
$\text{Criterion}(N_{i},N_{j})=\frac{f_{i}f_{j}}{||f_{i}||||f_{j}||}$ $-\text{MSE}(p_{i},p_{j})$ (1)
3.1.2 Dynamic Action Planning
As shown in Fig 2, we then input the topological memory $G_{t}$ at time $t$ into a multimodal transformer to obtain representations. this multimodal transformer contains pre-trained real node ( [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] ) encoders, fine-tuning imagination node ( [Uncaptioned image] ) encoders, and a pre-trained instruction encoder. The real node encoder and the imagination node encoder have the same structure and are therefore unified as node encoder for elaboration.
Node Embedding. The node visual feature $\hat{V}_{t}$ comprises three elements: node feature $f_{t}$, location encoding, and navigation step encoding. The latter marks visited nodes with their last visited time (00 for [Uncaptioned image], [Uncaptioned image]). This encoding reflects the temporal structure of memory. $\hat{V}_{t}$ is categorized into $\hat{V}_{t}^{r}$ and $\hat{V}_{t}^{i}$ based on whether the nodes are real or imaginary. Furthermore, a ‘stop’ node is added to the memory to represent a stop action and is connected to all other nodes.
Instruction and Node Encoders. Each word embedding in $I$ is augmented with a positional embedding and a type embedding. A multi-layer transformer processes these tokens to generate contextual representations, denoted as $\hat{I}$
Node and word embeddings are fed into a multi-layer cross-modal transformer. Inspired by (Chen et al. 2022), we endow it with a graph-aware self-attention (GASA) layer to capture the environment layout. The network output is navigation scores $s_{i}^{r}$ for navigable nodes and $s_{i}^{i}$ for imagination nodes. FFN below is a two-layer feed-forward network.
$s_{t}^{k}=\text{FFN}(\text{GASA}(\hat{V}_{t}^{k})),k=i,r$ (2)
Action Fusion Policy. Our work introduces an adaptive framework for navigation decision-making, which involves the dynamic integration of navigation scores obtained from both navigable nodes and imagination nodes. Initially, navigation scores from imagination nodes will be multiplied by a fusion factor and added to the scores of the nearest navigable nodes in Euclidean distance, since they are not reachable. The fusion factor is generated by concatenating the imagination nodes and real nodes and then feeding it into an FFN layer. The network structure is illustrated as follows.
$\gamma_{t}=\text{Sigmoid}(\text{FFN}(\left[\hat{V}^{r}_{t},\hat{V}^{i}_{t}\right]))$ (3)
$\hat{s_{t}}=s_{t}^{r}+\sum\nolimits_{j\in\mathcal{S}(i)}\gamma_{t}s^{i}_{t,j}$ (4)
where $\mathcal{S}(i)$ represents the set of all imagination nodes that need to be added for navigable node $N_{i}$.
3.2 Recurrent Imagination Tree
We put forth a recurrent local imagination tree mechanism. Drawing inspiration from (Koh et al. 2021; Mallya et al. 2020), the SALI agent can generate high-resolution future image information. For navigation memory $G_{t}$ at step $t$, a length-$K$ history information $H_{t}=\{d_{t-K:t},s_{t-K:t},p_{t-K:t}\}$, the RGB $r_{t}$, and the neighboring position of the current point $p^{g}_{t}$ are utilized to initialize imagery tree $T_{t}^{0}=\{H_{t},r_{t},p^{g}_{t}\}$. The imagery tree $T_{t}^{M}$ is then generated iteratively and finally integrated with the $G_{t}$ as mentioned in Section 3.1.1.
3.2.1 Recurrent Imagination Mechanism
The generation of the imagination tree is an iterative process. Without loss of generality, Fig 3 illustrates the process of expanding a recurrent imagination tree of $t+N+1$ ($N+1\leq M$) from $t+N$. We first use the input queue $T^{N}_{t}$ to generate a list of structured label images $s_{t+N+1}$ and $d_{t+N+1}$ at moment $t+N+1$ by the inpaint model for each navigable position $p^{g}_{t+N}$. Subsequently, the guidance image $g_{t+N}$ is generated through a point cloud projector. Finally, the RGB image $r_{t+N+1}$ and the target position $p^{g}_{t+N+1}$ will be generated by the spade model with $g_{t+N}$, $s_{t+N+1}$, and $d_{t+N+1}$ as inputs. The short-term memory is then updated to the state $T^{N+1}_{t}$. It is noteworthy that all images are uniformly oriented with 00 headings to prevent the generation of different images for the same point due to variations in agent action orientations. And we set the history length $K$ as 2 for the imagination part.
3.2.2 High-fidelity Image Generation
1) Inpaint Model. The inpaint model is an encoder-decoder-based network designed for structured label image generation. According to different $p^{g}_{t+N}$, the initial step will convert input data, represented by $H_{t}$, into corresponding structural maps $p_{t+N+1}$ through a point cloud projector. Subsequently, the map is fed into an encoder-decoder based on RedNet and ResNet (He et al. 2016) to generate one-hot semantic code $o_{t+N+1}$ and $d_{t+N+1}$. The structure of the inpaint model is as follows.
[…]
$[d_{t+1},o_{t+1}]$ $=\sigma C_{S}(X)+(1-\sigma)C_{F}(X_{t})$ (6)
$\sigma$ is denoted as weight parameter $C_{G}(X_{t})$. $C_{G},C_{S},C_{F}$ stand for Gate-, Shortcut- and Final-Convolution layer.
Figure 4: Before and after adding room-type model.
Nevertheless, the items in the generated semantic image will become fuzzy and illusory with higher $M$ (as illustrated in Fig 4) (Kaneko and Harada 2021). The likelihood of different items in the same pixel would be similar. For this reason, we incorporate the room type model. Given the high correlation between room information and items (e.g., refrigerator-kitchen, bed-bedroom), the SALI agent can ascertain the presence of specific items by determining the room type based on commonsense knowledge. We get the object weight dictionary $w$ for each room and revise the $o_{t+1}$ as $o_{t+1}\cdot(1+w)$. Then it’s transformed to get $s_{t+1}$.
2) Spade Model. The spade model is an image-to-image translation GAN (Wang et al. 2018a; Goodfellow et al. 2020) that enables high-resolution RGB generation capability (Park et al. 2019). The input to the spade model includes the embeddings of $r_{t+N}$ and $g_{t+N}$, and the output is the RGB image $r_{t+N+1}$.
For the following waypoint $p^{g}_{t+N+1}$ prediction, learned $r_{t+N+1}$ and $d_{t+N+1}$ are fed into a waypoint model, a prediction model based on bidirectional BERT.
3.2.3 Cross-correction
The mechanisms of the global memory and local imagination are mutually supportive. For the global memory, the imagination can assist in enhancing the image features of unvisited nodes ( [Uncaptioned image] ) (Section 3.1.1) and expanding the network structure. For the local imagination, the global map provides the necessary historical information and multiple trajectories, reducing information loss, for a certain imagination point.
3.3 Training and Inference
3.3.1 Multimodal Transformer Pre-training
Previous studies have shown the benefits of initializing VLN models with auxiliary tasks. Accordingly, we pre-train our model using expert behavior and imitation learning. The tasks are as follows: masked language modeling (MLM), masked region classification (MRC) (Lu et al. 2019), single-step action prediction (SAP) (Chen et al. 2021), and object grounding (OG) (Lin, Li, and Yu 2021).
Val Unseen-R2R Test Unseen-R2R Val Unseen-REVERIE Test Unseen-REVERIE
[…]
Seq2Seq (Anderson et al. 2018) 7.81 28 21 - 7.85 27 20 - 56 36 26 55 32 22
RecBert (Hong et al. 2021) 3.93 - 63 57 4.09 70 63 57 30 18 15 29 16 13
HOP+ (Qiao et al. 2023) 3.49 - 67 61 3.71 - 66 60 36 22 19 34 20 17
DUET (Chen et al. 2022) 3.31 81 72 60 3.65 76 69 59 46 32 23 52 31 22
BEVBert (An et al. 2023) 2.81 84 75 64 3.13 81 73 62 51 34 24 52 32 22
Lily (Lin et al. 2023) 2.48 84 77 72 3.05 82 74 68 48 32 23 45 30 21
ScaleVLN (Wang et al. 2023b) 2.34 87 79 70 2.73 83 77 68 57 - - 56 - -
SALI (Ours) 1.92 86 82 78 2.08 83 79 74 58 38 28 56 34 25
Table 1: Comparison with SoTA methods on R2R and REVERIE benchmarks.
3.3.2 Imagination Model pre-training
1) Inpaint Model. We trained the inpaint model to minimize losses of a cross-entropy loss and a mean absolute error (MAE) term. During the training process, random noise is added to the input.
$\mathcal{L}_{inpaint}=-\lambda\sum\nolimits_{i}\hat{s}_{i}\log(s_{t})+(1-\lambda)\|d_{t}-\hat{d}_{t}\|_{1}$ (7)
$\lambda$ is weight parameter. $s_{t},d_{t}$ are generated semantic- and depth-images, while $\hat{s}_{t},\hat{d}_{t}$ are ground-truth images.
For room-type model training, we get the room-type one-hot codes for each viewpoint from Matterport3D. Two ResNet-50 networks (Russakovsky et al. 2015), which have been previously trained on the ImageNet dataset and VLN data, are employed to encode the semantic image $s$ and the depth image $d$, respectively. The room-type model takes in the encoded feature and outputs one-hot codes. The loss is the cross-entropy between the predicted label $l$ and true room type label $\hat{l}$.
$\mathcal{L}_{room}=-\lambda\sum\nolimits_{i}\hat{l}_{i}\log(l_{i})$ (8)
2) Spade Model. The spade model is a GAN-based model. For the generator, we employ GAN hinge loss, feature matching loss, and perceptual loss (Johnson, Alahi, and Fei-Fei 2016). For PatchGAN-based discriminator (Phillip et al. 2017), the loss function is the loss for both ground-truth images $\hat{r}_{t}$ and generated images $r_{t}$. $\phi^{i}$ and $D^{i}$ denote the output of the $i$-th layer of the pre-trained VGG-19 model and discriminator.
$\mathcal{L}_{spade}^{d}=$ $-\mathbb{E}[\min(0,-1+D(\hat{r}_{t}))]$
$-\mathbb{E}[\min(0,-1+D(r_{t}))]$ (9)
$\mathcal{L}_{spade}^{g}=$ $-\lambda_{G}\mathbb{E}[D(r_{t})]$
$+\lambda_{F}\sum\nolimits_{i}^{n}\frac{\|\phi^{i}(\hat{r}_{t})-\phi^{i}(r_{t})\|_{1}}{n}$
$+\lambda_{P}\sum\nolimits_{i}^{n}\frac{\|D^{i}(\hat{r}_{t})-D^{i}(r_{t})\|_{1}}{n}$ (10)
The waypoint model is a BERT-based waypoint predictor (Hong et al. 2022). For each viewpoint, we labeled its neighboring points’ relative headings and distances into a 120 $\times$ 12 matrix representing 360 degrees and 3 meters (each element represents 3 degrees and 0.25 meters). A heat map, designated as $\hat{m}\in\mathbb{R}^{120\times 12}$, is then generated as the ground-truth data through interpolation of the aforementioned matrix. For the training process, the above-mentioned two ResNet-50 networks are employed to encode the RGB image and depth image for embedding, $e^{r}$ and $e^{d}$. Subsequently, $e^{r},e^{d}$ are merged by a non-linear layer and fed into the model, which generates a heatmap, $m\in\mathbb{R}^{120\times 12}$. Subsequently, non-maximum suppression (NMS) (Hosang, Benenson, and Schiele 2017) is applied to obtain neighboring waypoints. The loss is as follows.
$\mathcal{L}_{waypint}=\frac{1}{120\times 12}\sum\nolimits_{i=1}^{120}\sum\nolimits_{j=1}^{12}(m_{ij}-\hat{m}_{ij})^{2}$ (11)
3.3.3 Inference
At each stage of the testing process, the online agent constructs a global map of an imagination-based memory. It then reasons about the next action over the map, as explained in Section 3.1. As the memory expands, the SALI agent will approach the correct point and find the shortest path to the target. The agent will stop if it selects the ‘stop’ node or reaches the number limit of action steps. The best model checkpoint is selected by SPL$+$SR.
Figure 5: Comparison of SR and SPL between SALI and BEVBert models on different sub-datasets.
4 Experiments
4.1 Task Setup
Datasets. We evaluate SALI on VLN benchmarks with both fine-grained instructions (R2R) (Anderson et al. 2018) and high-level ones (REVERIE) (Qi et al. 2020). R2R provides step-by-step instructions with an average length of 32 words. REVERIE gives pre-defined object bounding boxes and instructions to guide agents in describing target objects’ positions. Instructions consist of 21 words on average.
Evaluation Metrics. We adopt the evaluation metrics used commonly by existing works (Qi et al. 2020): 1) Navigation Error (NE) calculates the distance between stop locations and target ones; 2) Trajectory Length (TL) represents the average path length; 3) Success Rate (SR) shows the ratio of stopping within 3m to the target;4) Success rate weighted by Path Length (SPL) makes trade-off between SR and TL; 5) Oracle Success Rate (OSR) is the ratio of including a viewpoint along the path where the target position can be seen; 6) Remote Grounding Success (RGS) is the ratio of successfully executed instructions; 7) RGSPL is RGS penalized by path length.
4.2 Comparison with State-of-the-art
R2R. Table 1 compares the navigation performance of the SALI agent with a range of approaches on R2R. The SALI agent outperforms the other approaches in all matrices. It is notable that SR increases concurrently with SPL, indicating that the SALI facilitates more correlation between scenes and instruction, thereby enhancing navigation efficiency.
REVERIE. SALI also achieves the best performance on the REVERIE benchmark. SALI surpasses the previous best model. We also note that the RGS and RGSPL increase a lot. We attribute it to the fact that the imagination model, especially the room-type model can help the agent recognize objects around itself better.
4.3 Quantitative and Qualitative Analysis
In navigation, the agent’s capacity for episodic memory and episodic simulation is reflected in its grasp of the environment-related commonsense. This is particularly evident in the agent’s ability to process complex associative information (e.g., room and object type associations, location associations between rooms).
Quantitative Study. We examine agents’ navigation ability under instructions with complex item and room information. We divided the R2R val_unseen instruction set into:
  • $\mathcal{S}_{1}$: Instructions have over 2 room words and only room words. (e.g. “Walk through hallway and towards restroom.” )
  • $\mathcal{S}_{2}$: Instructions have over 2 object words and only object words (e.g. “Turn left at the oven and past the fridge.” )
  • $\mathcal{S}_{3}$: Complex instructions have over 4 object and room words. (e.g. “Past the door next to the TV in the living room, then walk into the kitchen.” )
We then tested the SALI agent on the three instruction sets as illustrated in Fig 5. Compared to the previous best space-aware model (BEVBert, (An et al. 2023)), SALI wins the most (bandwidth represents a performance gap) on $\mathcal{S}_{3}$ in SR and SPL. SALI’s capabilities of imaging spatial information enable itself to effectively comprehend the interrelationship between objects and rooms, thus performing better when instruction comprises intricate item and room information. We demonstrated SALI achieves commonsense ability through our imagination-memory mechanism, which is analogous to the human episodic simulation memory.
Figure 6: Imagined picture display and waypoint prediction schematic. The imagination error is visualized by the curve.
Qualitative Analysis. We visualize the process of generating spatial and image information by the imagination mechanism in Fig 6. The left side shows the imagination for different distance goal positions and the corresponding real images. The peak signal-to-noise ratio (PSNR) (Korhonen and You 2012) of the generated image and the Pearson correlation coefficient (Cohen et al. 2009) with the real image were calculated. The curved area in the figure reflects the mean and variance of the pixels’ error of the two images, which can be approximated by visualizing the Pearson correlation coefficient and PSNR. For spatial waypoint prediction, the predicted heat map $m_{t}$ after the nms operation (top right) is visualized to show neighboring navigable positions.
4.4 Ablation Study
We conduct extensive experiments on the R2R val-unseen split to evaluate our key design choices.
Imagination vs. Reality. We first investigated the role of the imagined-memory hybrid mechanism in improving navigation performance, see Table 2. It indicates that the memory-based imagery mechanism gets the SoTA performance, which is 12% and 9% higher than “reality only” in SR and SPL. At the same time, we find that navigation using only imagination is the least effective. We attribute this to the inability of the agent to effectively use and correlate the imagination results with the navigation experience and trajectory. This suggests it’s necessary to establish long-term memory mechanisms supporting imaginative abilities.
[…]
1 Reality 12.59 3.39 78 70 61
2 Imagination 15.71 5.44 67 58 50
3 Reality + Imagination 10.34 1.92 86 82 70
Table 2: Ablation study of imagination-based memory.
Temporal and Spatial Imagery. We investigated the effect of the spatio-temporal range of memory on navigation ability. We use the navigation output time step $M$ and the upper limit of the number of nodes $\bar{N}$ to represent the memory range, see Table 3 (Input length is fixed: $K=2$). We can tell that the navigation ability increases with the expansion of the memory. If the navigation range is too large ($M=2,\bar{N}=8$), however, the SPL will decrease. This is because too large a range makes agents at different positions imagine the same destination, which interferes with the agent to make the optimal decision. Furthermore, it is important to consider the trade-off between navigation performance and training time per epoch. However, the overall inference time per epoch is relatively consistent and does not significantly impact the results.
[…]
1 0 0 78 70 61 0.54h 0.15h
2 1 4 82 76 67 0.74h 0.18h
3 2 4 86 82 71 1.32h 0.25h
4 2 8 84 82 68 2.51h 0.30h
Table 3: The effect of imagination range.
Auxiliary Model Effectiveness. The room type model and the waypoint model in the imagination model are auxiliary models that are not necessary steps for image generation. Waypoint predictions can be made by setting random directions instead. We examined the effect of the two models, see Table 4. We found the agent containing the auxiliary model is optimal in the SPL metrics, reflecting that the two models enable efficient decision-making.
[…]
1 None 12.59 3.39 78 70 61
2 Room 12.32 2.88 82 75 64
3 Waypoint 11.85 2.54 84 77 66
4 Room + Waypoint 10.34 1.92 86 80 71
Table 4: Ablation study of imagination auxiliary models.
Dynamic Decisioning. The effect of imagination and real node embeddings on the navigation decision is weighted. Comparative experiments are designed, one with dynamic weights (Section 3.2.2) and the other with fixed weights. We observed that dynamic weighting is more effective, with the imagination module’s weight decreasing as navigation progresses, which suggests the later stages of navigation rely less on imagination, akin to human memory mechanism.
[…]
1 Dynamic 10.34 1.92 86 80 71
2 Fixed ($\gamma_{t}=0.5$) 12.33 2.25 82 76 67
Table 5: Ablation study of Decision Weight.
5 Conclusion
The SALI agent is a VLN task agent with imagination-based memory, which is analogous to human episodic memory and episodic simulation. Extensive experimentation has demonstrated the credibility of the imagined results. Our agents have achieved SoTA performance in unseen environments. The future direction of work will focus on reducing the imagination time while maintaining fine-grained results and improving the generalization ability of the imagery module.